OpenStudy (anonymous):

Identify the variable that can be eliminated by adding the given equations in the given system of equations. ( probs. below)

OpenStudy (anonymous):

A- x + 7y=5, -x +2y=13 B- 2x-y=6, 3x + y =6 C- -2x + 5y= 10, 2x+3y =2

OpenStudy (ranga):

In A) if you add the two equations, the +x and -x will cancel out. So the variable x will be eliminated.

OpenStudy (anonymous):

I know how to do the problem, yet I don't understand what they mean by the eliminated variable. So, I would say for each problem the answer of what the variable's value is?

OpenStudy (ranga):

They are asking: If you add the two equations, one of the variable will disappear. They are asking which variable will disappear: x or y? They are not asking you to solve for x and y. They just want you to name which variable will get eliminated (or go away) when you add the two equations.